OpenStudy (anonymous):

How many solutions does this system of equations have? x-3y=6 9y=3x-18

OpenStudy (anonymous):

Infinite: 9y=3x-18 3y=x-6 x-3y=6

OpenStudy (anonymous):

you are correct

OpenStudy (anonymous):

What kind of system is the set of equations below? 5x - 4y = -3 -5x + 4y = 3

OpenStudy (anonymous):

What is the solution set of the given equation? 4y - 6y + 9y = -2